# Reinforcement Learning for Self-Improving Agent with Skill Library

### v1

@@ Abstract
Large Language Model (LLM)-based agents have demonstrated remarkable capabilities in complex reasoning and multi-turn interactions but struggle to continuously improve and adapt when deployed in new environments. One promising approach is implementing skill libraries that allow agents to learn, validate, and apply new skills. However, current skill library approaches rely primarily on LLM prompting, making consistent skill library implementation challenging. To overcome these challenges, we propose a Reinforcement Learning (RL)-based approach to enhance agents’ self-improvement capabilities with a skill library. Specifically, we introduce Skill Augmented GRPO for self-Evolution (SAGE), a novel RL framework that systematically incorporates skills into learning. The framework’s key component, Sequential Rollout, iteratively deploys agents across a chain of similar tasks for each rollout. As agents navigate through the task chain, skills generated from previous tasks accumulate in the library and become available for subsequent tasks. Additionally, the framework enhances skill generation and utilization through a Skill-integrated Reward that complements the original outcome-based rewards. Experimental results on AppWorld demonstrate that SAGE, when applied to supervised-finetuned model with expert experience, achieves 8.9% higher Scenario Goal Completion while requiring 26% fewer interaction steps and generating 59% fewer tokens, substantially outperforming existing approaches in both accuracy and efficiency.
@@ 1 Introduction
Large language model (LLM)-based agents have been widely applied to automate complex tasks through active environmental interactions, including coding agent (yang2024swe; novikov2025alphaevolve), deep research (deep_research), assistant agent yao2024tau; chen2025reinforcement, and web browsing (yao2022webshop; zhouwebarena). To enhance the performance of these multi-turn interactive agents, researchers have successfully integrated reinforcement learning (RL) techniques into their frameworks (chen2025reinforcement; qiwebrl; zhou2025sweet; wang2025ragen). Recent advances, particularly in reinforcement learning with verifiable rewards (RLVR) (shao2024deepseekmath; guo2025deepseek), have enabled effective end-to-end agent training for improved performance (jin2025search). However, despite RL’s effectiveness, significant limitations persist: RL-trained agents are often limited to specific training scenarios (zheng2024gpt; li2024effects). When deployed in new environments, they struggle to demonstrate continual learning capabilities to effectively utilize valuable on-going experiences for future tasks.
@@
Recent research has made significant strides in enabling agents to compose reusable skills (wangvoyager; cailarge; nguyen2024dynasaur; wang2024agent; zheng2025skillweaver; wang2025inducing). For instance, wang2025inducing focuses on inducing high-level web browsing skills from successful action trajectories, transforming primitive actions like click and search into more complex operations such as search product. While these works have demonstrated the effectiveness of skill libraries for deployed agents, they predominantly rely on manually crafted prompts for skill generation and utilization. This prompt-based approach, however, is inherently constrained by the instruction following capabilities of the base model, limiting both the quality and adaptability of the skill library.
@@
In this paper, we explore enhancing agents’ self-improvement capabilities through RL with a skill library. Given the diverse frameworks of agents for various tasks, we focus specifically on tool-using agents that interact with environments through API calls and receive corresponding feedback. Our implementation extends the CodeAct framework (wang2024executable), which enables agents to compose code by combining multiple APIs with basic programming constructs (such as for loops) to solve complex tasks. Building upon this foundation, we develop a specialized framework for self-improving agents with skill library (referred to as skill library agents for simplicity). Unlike previous frameworks such as Agent Skill Induction (wang2025inducing), which typically define reusable skills only after task completion, we implement a unified format for both task solving and skill generation, following nguyen2024dynasaur. In our approach, when the agent model interacts with the environment through APIs, it generates programmatic functions that can be saved as skills and subsequently called to execute, rather than using multiple APIs directly. Given that in-context prompting struggles to adapt open-source models to this newly proposed skill library agent, supervised fine-tuning (SFT) is first applied before RL using high-quality trajectories collected through expert experience generated by advanced LLMs.
@@
Based on the SFT model, we propose a novel RL framework for skill library agents. Traditional RL approaches typically consider rewards only for individual examples, limiting their scope to ongoing task performance. For skill library agents, the focus is on developing high-quality, reusable skills while improving accurate skill usage from the library for task optimization. To realize this, we extend Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) and propose Skill Augmented GRPO for self-Evolution (SAGE), consisting of Sequential Rollout and Skill-integrated Reward. Instead of single-task trajectories, the agent is trained with chains of similar tasks. We implement the Sequential Rollout process on the task chain, where skills generated in the previous tasks are preserved and made available for use in the following ones. Under this framework, the final Skill-integrated Reward is computed as the sum of two components: the verifiable outcome-based reward and an extra reward for high-quality skill generation and utilization.
@@
To evaluate the effectiveness of SAGE for skill library agents, we conducted experiments on the AppWorld dataset (trivedi2024appworld), where agents interact with their environment through API documentation lookup, API calls, and logical programming constructs to solve complex practical tasks. AppWorld utilizes the Scenario Goal Completion (SGC) metric, which measures the success rate of scenarios where all three similar tasks within a scenario are successfully completed. This metric effectively evaluates how well generated skills transfer across similar tasks. After applying SAGE to Qwen2.5-32B-Instruct, we observe significant improvements compared to prompting-based approaches. Our method achieves more than 3x SGC score while requiring less than half of generated tokens. Analysis of skill utilization reveals that agents trained with SAGE demonstrate more than 2x success rates when utilizing learned skills.
@@ 2 Related Work
LLM-based Agent. Recent advancements in instruction-tuned LLMs have enabled them to follow user instructions for interacting with external environments as autonomous agents. Various frameworks have been developed to enhance these backbone LLMs’ capabilities in performing agentic tasks. yao2023react pioneered a "reason-then-act" pipeline, guiding LLMs to generate interactive actions for agents after a text reasoning process. erdoganplan extended this approach by incorporating an additional planning phase. Additionally, wang2024executable demonstrated that generating executable Python code and then run it within a code interpreter could significantly improve agent performance. To further enhance agent capabilities, researchers have applied both supervised fine-tuning (schick2023toolformer; chen2023fireact; zeng2024agenttuning; zhang2024agentohana) and reinforcement learning (song2024trial; bai2024digirl; wang2025ragen; chen2025reinforcement; qiwebrl; zhou2025sweet) to develop more effective and adaptive agent backbone LLMs.
Self-Improving Agent with Skill Library. While RL algorithms have enabled agents to self-improve through valuable experiences explored during rollouts (zhouproposer; putta2024agent; qiwebrl), enabling continuous self-improvement after deployment, especially in new environments, remains challenging. wangvoyager pioneered the use of a skill library to record successful behaviors for later retrieval in Minecraft exploration. Subsequently, numerous studies have demonstrated the effectiveness of such skill libraries across various agentic tasks, including web exploration (wang2024agent; zheng2025skillweaver; wang2025inducing), computer control (zhengsynapse; wu2024copilot), and math problems (nguyen2024dynasaur). These skills can take two forms: natural language experience memories serving as a reference, or executable skills that can be directly implemented in the environment. In this paper, we focus on leveraging RL to enhance agents’ self-improvement capabilities by teaching them to generate executable skills for the skill library during test time.
@@ 3.1 Skill Library Agent
Existing frameworks for skill library agents, such as Agent Skill Induction (wang2025inducing), Agent Workflow Memory (wang2024agent), and Voyager (wangvoyager), typically define reusable skills after completing entire tasks. While this approach allows agents to observe complete task trajectories before determining skill definitions, it limits the RL process in two ways: (1) In long-horizon tasks, the additional skill generation process further extends the context length, potentially exceeding the model’s limitations; (2) The separation between task execution and skill generation creates an inconsistency that may impact learning effectiveness.
To address these limitations, we follow the DynaSaur approach nguyen2024dynasaur and implement a unified format for both task solving and skill generation. Specifically, when the agent model interacts with the API environments, it first generates a skill function and then calls it to process, instead of using multiple APIs. An example of the format difference is presented in Appendix A.
@@ 3.2.1 Preliminary
In this paper, we build our SAGE based on GRPO (shao2024deepseekmath). For each query $q$, GRPO first samples a group of outputs $\{o_{1},...,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy by maximizing the following objective:
@@ 3.2.4 SAGE
After collecting the rollout trajectories and computing the corresponding rewards, we implement our Skill Augmented GRPO for self-Evolution (SAGE) to train the skill library agent. Following chen2025reinforcement, we choose not to use the KL divergence penalty and the advantage would not be normalized by standard deviation of the rewards.
@@ 4.1 Experimental Settings
Dataset and Base Model. To evaluate the effectiveness of SAGE for skill library agents, particularly for those designed for long-horizon tool usage, we adopt the AppWorld dataset (trivedi2024appworld). AppWorld contains 750 tasks from 250 task scenarios. Each scenario consists of three tasks sharing similar instructions. The scenario-based structure of AppWorld makes it well-suited for Sequential Rollout, as tasks within the same scenario naturally form a task chain. We utilize the Train set for all training steps, including SFT and subsequent SAGE. The Dev set guides the best checkpoint selection during training. Evaluation is performed on both Test-Normal and Test-Challenge splits. More details about the dataset are presented in Appendix C.
To ensure a fair comparison with the previous work (chen2025reinforcement) on AppWorld, we used Qwen2.5-32B-Instruct (qwen2.5) as our base model for training purpose.
Skill Library Agent. Following the ReAct agent (yao2023react) as presented in the AppWorld paper (trivedi2024appworld), we implemented our skill library agent with a specifically designed in-context example and detailed instructions to guide the base model in using the skill library to perform complex tasks. The detailed prompt of our skill library agent is presented in Appendix D. Regarding skill retrieval, we employed an idealized case during sequential rollout where skills are retained and utilized only within the same scenario in the task chain. This approach eliminates the need for a retrieval model, significantly simplifying the RL rollout processes. While our default evaluation uses the ideal case (Same Scenario) with provided scenario labels from the test set, we conducted additional analysis on various retrieval methods to address real-world scenarios where such labels may not be available.
Baseline GRPO. Since the baseline method LOOP (chen2025reinforcement) does not provide source code or model, and their paper does not report agent efficiency metrics, it is challenging to make a direct, fair comparison. Therefore, we implemented our own baseline training using the GRPO method without Skill Library for comparison purposes. Following chen2025reinforcement, our baseline GRPO removes the KL divergence penalty and calculates advantage using the mean reward within the group instead of normalizing by the standard error. The only difference between our baseline GRPO and LOOP is that we perform strictly on-policy RL without PPO epochs. More training settings of Baseline GRPO are presented in Appendix E.
@@
Table 1: Task Performance of SAGE compared with various baselines. Results of Methods marked with "*" are taken from (chen2025reinforcement). In these cases, Avg. Steps and Tokens were not reported, thus denoted by "- -".
@@ Appendix C AppWorld Dataset
In AppWorld dataset (trivedi2024appworld), autonomous agents are required to complete everyday digital tasks (e.g., sending messages and transfering money to roommates) by consulting API documentation and executing API calls through generated code. Unlike other API usage datasets, AppWorld provides a high-quality execution environment that simulates 9 everyday apps (Amazon, Spotify, Venmo, Gmail, Todoist, SimpleNote, Splitwise, FileSystem, Phone, and ApiDocs). The environment includes totally 457 APIs and features interactions with over 100 simulated users.
@@
For accurate evaluation, each task incorporates a manually written program that evaluates the final environment state against predefined criteria, producing a completion rate (0-1) that serves as our outcome-based reward. The stage-based tests accommodate multiple solution paths, enabling more precise reward calculations. Notably, AppWorld presents a significant challenge even for state-of-the-art LLMs. According to trivedi2024appworld, even GPT-4O achieves only about 49% success rate on Test-Normal tasks and approximately 30% on Test-Challenge tasks.
@@ Appendix E Training Details of Baseline GRPO
We built our training framework based on the verl repository (sheng2024hybridflow), which uses vLLM (kwon2023efficient) for rollout and FSDP (zhao2023pytorch) for gradient update. While our training configuration largely followed chen2025reinforcement, there are still some key modifications: we implemented strictly on-policy reinforcement learning and employed full-parameter fine-tuning instead of LoRA. The implementation details of the baseline GRPO used in our paper are shown as follows:
@@
Training Settings. We performed the whole RL training process on 4xNVIDIA H100 8-GPU nodes. Due to the long-horizon trajectories, we utilized the ulysses sequence parallel (jacobs2023deepspeed), allocating each example on 4 GPUs. We kept the mini batch size the same as the 288 batch size of rollout. A constant learning rate of 1e-6 and clip the gradient norm to 1 were applied for policy gradient update.

### v2

@@ Abstract
Large Language Model (LLM)-based agents have demonstrated remarkable capabilities in complex reasoning and multi-turn interactions but struggle to continuously improve and adapt when deployed in new environments. One promising approach is implementing skill libraries that allow agents to learn, validate, and apply new skills. However, current skill library approaches rely primarily on LLM prompting, making consistent skill library implementation challenging. To overcome these challenges, we propose a Reinforcement Learning (RL)-based approach to enhance agents’ self-improvement capabilities with a skill library. Specifically, we introduce Skill Augmented GRPO for self-Evolution (SAGE), a novel RL framework that systematically incorporates skills into learning. The framework’s key component, Sequential Rollout, iteratively deploys agents across a chain of similar tasks for each rollout. As agents navigate through the task chain, skills generated from previous tasks accumulate in the library and become available for subsequent tasks. Additionally, the framework enhances skill generation and utilization through a Skill-integrated Reward that complements the original outcome-based rewards. Experimental results on AppWorld demonstrate that SAGE, when applied to supervised-finetuned model with expert experience, achieves 8.9% higher Scenario Goal Completion while requiring 26% fewer interaction steps and generating 59% fewer tokens, substantially outperforming existing approaches in both accuracy and efficiency. Our code is available at https://github.com/amazon-science/SAGE.
@@ 1 Introduction
Large language model (LLM)-based agents have been widely applied to automate complex tasks through active environmental interactions, including coding agent (Yang et al., 2024; Novikov et al., 2025), deep research (OpenAI, 2025), assistant agent Yao et al. (2024); Chen et al. (2025), and web browsing (Yao et al., 2022; Zhou et al., 2024). To enhance the performance of these multi-turn interactive agents, researchers have successfully integrated reinforcement learning (RL) techniques into their frameworks (Chen et al., 2025; Qi et al., 2025; Zhou et al., 2025a; Wang et al., 2025a). Recent advances, particularly in reinforcement learning with verifiable rewards (RLVR) (Shao et al., 2024; Guo et al., 2025), have enabled effective end-to-end agent training for improved performance (Jin et al., 2025). However, despite RL’s effectiveness, significant limitations persist: RL-trained agents are often limited to specific training scenarios (Zheng et al., 2024a; Li et al., 2024). When deployed in new environments, they struggle to demonstrate continual learning capabilities to effectively utilize valuable on-going experiences for future tasks.
@@
Recent research has made significant strides in enabling agents to compose reusable skills (Wang et al., 2024a; Cai et al., 2024; Nguyen et al., 2024; Wang et al., 2024c; Zheng et al., 2025; Wang et al., 2025b). For instance, Wang et al. (2025b) focuses on inducing high-level web browsing skills from successful action trajectories, transforming primitive actions like click and search into more complex operations such as search product. While these works have demonstrated the effectiveness of skill libraries for deployed agents, they predominantly rely on manually crafted prompts for skill generation and utilization. This prompt-based approach, however, is inherently constrained by the instruction following capabilities of the base model, limiting both the quality and adaptability of the skill library.
@@
In this paper, we explore enhancing agents’ self-improvement capabilities through RL with a skill library. Given the diverse frameworks of agents for various tasks, we focus specifically on tool-using agents that interact with environments through API calls and receive corresponding feedback. Our implementation extends the CodeAct framework (Wang et al., 2024b), which enables agents to compose code by combining multiple APIs with basic programming constructs (such as for loops) to solve complex tasks. Building upon this foundation, we develop a specialized framework for self-improving agents with skill library (referred to as skill library agents for simplicity). Unlike previous frameworks such as Agent Skill Induction (Wang et al., 2025b), which typically define reusable skills only after task completion, we implement a unified format for both task solving and skill generation, following Nguyen et al. (2024). In our approach, when the agent model interacts with the environment through APIs, it generates programmatic functions that can be saved as skills and subsequently called to execute, rather than using multiple APIs directly. Given that in-context prompting struggles to adapt open-source models to this newly proposed skill library agent, supervised fine-tuning (SFT) is first applied before RL using high-quality trajectories collected through expert experience generated by advanced LLMs.
@@
Based on the SFT model, we propose a novel RL framework for skill library agents. Traditional RL approaches typically consider rewards only for individual examples, limiting their scope to ongoing task performance. For skill library agents, the focus is on developing high-quality, reusable skills while improving accurate skill usage from the library for task optimization. To realize this, we extend Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and propose Skill Augmented GRPO for self-Evolution (SAGE), consisting of Sequential Rollout and Skill-integrated Reward. Instead of single-task trajectories, the agent is trained with chains of similar tasks. We implement the Sequential Rollout process on the task chain, where skills generated in the previous tasks are preserved and made available for use in the following ones. Under this framework, the final Skill-integrated Reward is computed as the sum of two components: the verifiable outcome-based reward and an extra reward for high-quality skill generation and utilization.
@@
To evaluate the effectiveness of SAGE for skill library agents, we conducted experiments on the AppWorld dataset (Trivedi et al., 2024), where agents interact with their environment through API documentation lookup, API calls, and logical programming constructs to solve complex practical tasks. AppWorld utilizes the Scenario Goal Completion (SGC) metric, which measures the success rate of scenarios where all three similar tasks within a scenario are successfully completed. This metric effectively evaluates how well generated skills transfer across similar tasks. After applying SAGE to Qwen2.5-32B-Instruct, we observe significant improvements compared to prompting-based approaches. Our method achieves more than 3x SGC score while requiring less than half of generated tokens. Analysis of skill utilization reveals that agents trained with SAGE demonstrate more than 2x success rates when utilizing learned skills.
@@ 2 Related Work
LLM-based Agent. Recent advancements in instruction-tuned LLMs have enabled them to follow user instructions for interacting with external environments as autonomous agents. Various frameworks have been developed to enhance these backbone LLMs’ capabilities in performing agentic tasks. Yao et al. (2023) pioneered a "reason-then-act" pipeline, guiding LLMs to generate interactive actions for agents after a text reasoning process. Erdogan et al. (2025) extended this approach by incorporating an additional planning phase. Additionally, Wang et al. (2024b) demonstrated that generating executable Python code and then run it within a code interpreter could significantly improve agent performance. To further enhance agent capabilities, researchers have applied both supervised fine-tuning (Schick et al., 2023; Chen et al., 2023; Zeng et al., 2024; Zhang et al., 2024) and reinforcement learning (Song et al., 2024; Bai et al., 2024; Wang et al., 2025a; Chen et al., 2025; Qi et al., 2025; Zhou et al., 2025a) to develop more effective and adaptive agent backbone LLMs.
Self-Improving Agent with Skill Library. While RL algorithms have enabled agents to self-improve through valuable experiences explored during rollouts (Zhou et al., 2025b; Putta et al., 2024; Qi et al., 2025), enabling continuous self-improvement after deployment, especially in new environments, remains challenging. Wang et al. (2024a) pioneered the use of a skill library to record successful behaviors for later retrieval in Minecraft exploration. Subsequently, numerous studies have demonstrated the effectiveness of such skill libraries across various agentic tasks, including web exploration (Wang et al., 2024c; Zheng et al., 2025; Wang et al., 2025b), computer control (Zheng et al., 2024b; Wu et al., 2024), and math problems (Nguyen et al., 2024). These skills can take two forms: natural language experience memories serving as a reference, or executable skills that can be directly implemented in the environment. In this paper, we focus on leveraging RL to enhance agents’ self-improvement capabilities by teaching them to generate executable skills for the skill library during test time.
@@ 3.1 Skill Library Agent
Existing frameworks for skill library agents, such as Agent Skill Induction (Wang et al., 2025b), Agent Workflow Memory (Wang et al., 2024c), and Voyager (Wang et al., 2024a), typically define reusable skills after completing entire tasks. While this approach allows agents to observe complete task trajectories before determining skill definitions, it limits the RL process in two ways: (1) In long-horizon tasks, the additional skill generation process further extends the context length, potentially exceeding the model’s limitations; (2) The separation between task execution and skill generation creates an inconsistency that may impact learning effectiveness.
To address these limitations, we follow the DynaSaur approach Nguyen et al. (2024) and implement a unified format for both task solving and skill generation. Specifically, when the agent model interacts with the API environments, it first generates a skill function and then calls it to process, instead of using multiple APIs. An example of the format difference is presented in Appendix A.
@@ 3.2.1 Preliminary
In this paper, we build our SAGE based on GRPO (Shao et al., 2024). For each query $q$, GRPO first samples a group of outputs $\{o_{1},...,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy by maximizing the following objective:
@@ 3.2.4 SAGE
After collecting the rollout trajectories and computing the corresponding rewards, we implement our Skill Augmented GRPO for self-Evolution (SAGE) to train the skill library agent. Following Chen et al. (2025), we choose not to use the KL divergence penalty and the advantage would not be normalized by standard deviation of the rewards.
@@ 4.1 Experimental Settings
Dataset and Base Model. To evaluate the effectiveness of SAGE for skill library agents, particularly for those designed for long-horizon tool usage, we adopt the AppWorld dataset (Trivedi et al., 2024). AppWorld contains 750 tasks from 250 task scenarios. Each scenario consists of three tasks sharing similar instructions. The scenario-based structure of AppWorld makes it well-suited for Sequential Rollout, as tasks within the same scenario naturally form a task chain. We utilize the Train set for all training steps, including SFT and subsequent SAGE. The Dev set guides the best checkpoint selection during training. Evaluation is performed on both Test-Normal and Test-Challenge splits. More details about the dataset are presented in Appendix C.
To ensure a fair comparison with the previous work (Chen et al., 2025) on AppWorld, we used Qwen2.5-32B-Instruct (Qwen, 2024) as our base model for training purpose.
Skill Library Agent. Following the ReAct agent (Yao et al., 2023) as presented in the AppWorld paper (Trivedi et al., 2024), we implemented our skill library agent with a specifically designed in-context example and detailed instructions to guide the base model in using the skill library to perform complex tasks. The detailed prompt of our skill library agent is presented in Appendix D. Regarding skill retrieval, we employed an idealized case during sequential rollout where skills are retained and utilized only within the same scenario in the task chain. This approach eliminates the need for a retrieval model, significantly simplifying the RL rollout processes. While our default evaluation uses the ideal case (Same Scenario) with provided scenario labels from the test set, we conducted additional analysis on various retrieval methods to address real-world scenarios where such labels may not be available.
Baseline GRPO. Since the baseline method LOOP (Chen et al., 2025) does not provide source code or model, and their paper does not report agent efficiency metrics, it is challenging to make a direct, fair comparison. Therefore, we implemented our own baseline training using the GRPO method without Skill Library for comparison purposes. Following Chen et al. (2025), our baseline GRPO removes the KL divergence penalty and calculates advantage using the mean reward within the group instead of normalizing by the standard error. The only difference between our baseline GRPO and LOOP is that we perform strictly on-policy RL without PPO epochs. More training settings of Baseline GRPO are presented in Appendix E.
@@
Table 1: Task Performance of SAGE compared with various baselines. Results of Methods marked with "*" are taken from (Chen et al., 2025). In these cases, Avg. Steps and Tokens were not reported, thus denoted by "- -".
@@ Appendix C AppWorld Dataset
In AppWorld dataset (Trivedi et al., 2024), autonomous agents are required to complete everyday digital tasks (e.g., sending messages and transfering money to roommates) by consulting API documentation and executing API calls through generated code. Unlike other API usage datasets, AppWorld provides a high-quality execution environment that simulates 9 everyday apps (Amazon, Spotify, Venmo, Gmail, Todoist, SimpleNote, Splitwise, FileSystem, Phone, and ApiDocs). The environment includes totally 457 APIs and features interactions with over 100 simulated users.
@@
For accurate evaluation, each task incorporates a manually written program that evaluates the final environment state against predefined criteria, producing a completion rate (0-1) that serves as our outcome-based reward. The stage-based tests accommodate multiple solution paths, enabling more precise reward calculations. Notably, AppWorld presents a significant challenge even for state-of-the-art LLMs. According to Trivedi et al. (2024), even GPT-4O achieves only about 49% success rate on Test-Normal tasks and approximately 30% on Test-Challenge tasks.
@@ Appendix E Training Details of Baseline GRPO
We built our training framework based on the verl repository (Sheng et al., 2024), which uses vLLM (Kwon et al., 2023) for rollout and FSDP (Zhao et al., 2023) for gradient update. While our training configuration largely followed Chen et al. (2025), there are still some key modifications: we implemented strictly on-policy reinforcement learning and employed full-parameter fine-tuning instead of LoRA. The implementation details of the baseline GRPO used in our paper are shown as follows:
@@
Training Settings. We performed the whole RL training process on 4xNVIDIA H100 8-GPU nodes. Due to the long-horizon trajectories, we utilized the ulysses sequence parallel (Jacobs et al., 2023), allocating each example on 4 GPUs. We kept the mini batch size the same as the 288 batch size of rollout. A constant learning rate of 1e-6 and clip the gradient norm to 1 were applied for policy gradient update.